Fermi National Laboratory


Accelerator Update for April 29, 2005 - May 2, 2005

Friday April 29
The day shift began with Operations monitoring the Recycler stash, with the Tevatron (TeV) in shot setup, and with MiniBooNE in access.

Operators prepared the Meson enclosures M02 and M03 for access at 8:53 AM. Later, they prepared M04 for access.

Operations established combination store 4118 at 9:53 AM with an initial luminosity of 122.65E30. This store set a NEW LUMINOSITY RECORD.

At 10:16 AM, mechanical technicians reported a broken flag on a MiniBooNE magnet. This holds off MiniBooNE beam until it's repaired.

A Booster expert began a two-hour crawling wire study at 1:30 PM.

At 2:27 PM, a Booster RF station (BRF16) tripped off for the first of two times in eighteen minutes. It reset okay both times.

Operations put the Recycler into shot setup at 3:13 PM to stash antiprotons.

Recycler ended its stash at 4:08 PM.

The Antiproton Source (Pbar) began stacking at 4:31 PM.

Operations established beam to MiniBooNE at 4:30 PM.

At 4:59 PM, Operations reported a LCW leak alarm. Operators investigated and found a TeV sector F3 power supply leak (S:HP3US). Operators valved out the leak. An expert arrived and investigated further.

At 6:12 PM, a Booster power supply (ORBMP) tripped off. Operators replaced a blown fuse.

Experts reported at 7:48 PM that they had completed their repairs on HP3US. Besides fixing the LCW leak, experts also replaced four SCR modules and fixed a short.

Operations established beam to the Meson MCenter experiment at 8:18 PM.

At 7:56 PM, Operations reduced the Pbar stacking rate due to a problem with diagnostics.

At 9:09 PM, BRF16 tripped off. It reset okay.

A Pbar expert reported at 9:26 that he had repaired AP1001. Stacking resumed at its normal rate.

The Recycler began stashing antiprotons at 10:51 PM. The stashing ended and Pbar resumed stacking at 11:50 PM.

Saturday April 30
The midnight shift began with Operations monitoring store 4118 and the stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 12:59 AM, BRF15 tripped off. It reset okay.

At 1:05 AM, a Main Injector RF station (MIRF11) tripped off. It reset okay.

At 1:56 AM, a Linac RF station (LRF4) tripped off. It reset okay.

At 4:19 AM, BRF16 tripped off for the first of two times in fifteen minutes. It reset okay both times.

At 7:20 AM, a Linac RF station (KRF1) tripped off. Operators investigated and found a LCW flow fault. Operators had the station back on by 7:31 AM. KRF2 tripped off while operators were working on KRF1. KRF2 reset okay.

At 10:06 AM, Pbar had an emittance blowup. Operators adjusted an Accumulator RF station (ARF2), which helped.

At 10:23 AM, LRF1 tripped off. It reset itself.

At 11:32 AM, the Booster East Anode power supply (EAPS) tripped off with no status. Operators reset the station, but then it tripped off again. This happened several more times. Operators turned off the supply and called in an expert.

Operations terminated store 4118 around 12:58 AM. Operators made D-Zero (D0) ready for experimenters to access their collision hall.

At 1:13 PM, Operation had to halt stacking due to the Pbar emittance problems. Operations called in an expert who fixed the problem. Stacking resumed at 2:25 PM.

Operations established beam to NuMI for study purposes at 2:26 PM.

D0 came out of access at 3:01 PM.

At 3:44, a Booster vacuum pump tripped off and took BRF7 & 8 down with it. Everything reset okay.

Operations put the TeV into shot setup at 4:17 PM.

At 5:22 PM, an Accumulator water skid at AP30 tripped off. Operators investigated and then called in an expert. This problem put shot setup on hold.

Operations took the TeV out of shot setup and reestablished beam to MCenter at 7:09 PM.

Technicians had the Accumulator water skid replaced by 8:41 PM.

The TeV was back in shot setup by 9:21 PM.

Operations established store 4121 at 11:04 PM with an initial luminosity of 122.12E30.

MCenter resumed taking beam at 11:28 PM.

Pbar resumed stacking at 11:56 PM.

Sunday May 1
The midnight shift began with operations monitoring store 4121, with Pbar stacking, and with NuMI, MiniBooNE, and MCenter taking beam.

At 3:49 AM, the NuMI beam permit tripped off due to a beam problem. It reset okay.

The Recycler began stashing antiprotons at 6:31 AM.

At 7:25 AM, a Booster beam problem required operators to halt beam.

The Recycler stash ended at 7:38 AM.

Booster beam resumed 7:46 AM. MCenter beam and NuMI studies resumed. Pbar also resumed stacking.

At 7:49 AM, the NuMI beam permit tripped off for the first of two times in three minutes. It reset okay.

At 9:49 AM, BRF16 tripped off. It reset okay.

NuMI experts reported at 10:40 AM that they had completed their target scan. They requested an intensity increase. This caused some MI slipstacking beam problems.

At 11:44 AM, BRF1 tripped off. It reset okay.

At 1:04 PM, Pbar suffered another emittance blowup. The problem resolved itself just as operators began tuning.

The Recycler began stashing antiprotons at 3:22 PM and finished about an hour later.

At 3:24 PM, the Booster EAPS tripped off. It reset okay.

Pbar resumed stacking at 4:57 PM,

MCenter resumed taking beam at 5:13 PM.

Two MI experts reported at 8:07 PM that they had resolved MI beam problems.

At 10:41 PM, BRF16 tripped off. It reset okay.

At 11:33 PM, NuMI studiers tried to turn on a kicker power supply (KPS6N), but it tripped off.

At 11:38 PM, the NuMI beam permit tripped off several times due to a device trip (VP609N) trip. At least one of the trips was due to a studier. It reset okay.

Monday May 2
The midnight shift began with Operations monitoring store 4121 and the stash, with Pbar stacking, with NuMI experts conducting studies, and with MiniBooNE and MCenter taking beam.

At 12:51 AM, the NuMI beam permit tripped off due to a beam problem. The studiers asked to hold beam off until they could find the problem. At 2:04 AM they reported they had found a reading that suggested a helium leak. They called for expert help.

NuMI experts requested an access to their target hall at 3:13 AM. Operations called in a safety expert. Operators prepared NuMI for access at 4:17 AM.

Operations terminated store 4121 at 5 AM and put the TeV into shot setup fifteen minutes later.

Operations established store 4122 at 7:24 AM with an initial luminosity of 122.13E30. Pbar resumed stacking and MCenter resumed taking beam.

The Plan for Monday
The plan for today is to monitor the store and to continue stacking and stashing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/2/2005   email Fermilab